1. A semicircular window of radius 15 inches is to be laminated with a sunblock coating that costs $0.85 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use p = 3.14. Hint: Area of Circle: A = prē)
Answer by rfer(10417)

A=pir^2/2
A=3.14(15)^2/2
A=3.14(225)/2
A=353.25 sq in *0.85=$300.26